Michael purchased a Wayne Gretzkey rookie card for $3.50 in 1985. In 2008 the card is worth $58.19. What is the annual rate of appreciation??? I think it is 1.3%. Can someone please help me ???
Answer by scott8148(6628)

your decimal point is in the wrong place

(present value)=(starting value)(1+((percent appreciation)/100))^(time)

let x="%/100" __ 58.19=3.5(1+x)^23

dividing by 3.5 and taking log __ log(58.19/3.5)=(23)log(1+x)

dividing by 23 __ (log(58.19/3.5))/23=log(1+x)

taking antilog and subtracting 1 __ 10^((log(58.19/3.5))/23)-1=x